OpenStudy (anonymous):

Transform the polar equation to an equation in rectangular coordinates: r = 8sin(Theta). Write an equation in rectangular coordinates.

OpenStudy (freckles):

use r^2=x^2+y^2 and r sin(theta)=y and r cos(theta)=x and tan(theta)=y/x hint: you will only need two of these

OpenStudy (freckles):

for this one

OpenStudy (freckles):

another free hint: multiply both sides by r